How many atoms do mercury vapour molecules consist of if the density of mercury vapour relative to air is 6.92? The average mass of air is 29g/mol

Grade:12

1 Answers

Sunil Kumar FP
9 years ago
(density of mercury vapour/density of air) =6.92
In a given volume
mass of mercury vapour/mass of air) =6.92
mass of mercury vapour =6.92*29 g/mole
=200.68 g/mole
therefore number of mercury molecule =Na*200.68 atoms
Na is the avagadro number
=6.023*10^23*200.68 atoms
=1.208*10^26 atoms